# Xiangqi Wang1, Xiuying Chen2, Mohamed Elhoseiny3, Xiangliang Zhang1,†

### v1

@@ Abstract
Large Vision-Language Models (LVLMs) have become essential for advancing the integration of visual and linguistic information, facilitating a wide range of complex applications and tasks. However, the evaluation of LVLMs presents significant challenges as the evaluation benchmark always demands lots of human cost for its construction, and remains static, lacking flexibility once constructed. Even though automatic evaluation has been explored in textual modality, the visual modality remains under-explored. As a result, in this work, we address a question: “Can LVLMs serve as a path to automatic benchmarking?”. We introduce AutoBench-V, an automated framework for serving evaluation on demand, i.e., benchmarking LVLMs based on specific aspects of model capability. Upon receiving an evaluation capability, AutoBench-V leverages text-to-image models to generate relevant image samples and then utilizes LVLMs to orchestrate visual question-answering (VQA) tasks, completing the evaluation process efficiently and flexibly. Through an extensive evaluation of seven popular LVLMs across five demanded user inputs (i.e., evaluation capabilities), the framework shows effectiveness and reliability. We observe the following: (1) Our constructed benchmark accurately reflects varying task difficulties; (2) As task difficulty rises, the performance gap between models widens; (3) While models exhibit strong performance in abstract level understanding, they underperform in details reasoning tasks; and (4) Constructing a dataset with varying levels of difficulties is critical for a comprehensive and exhaustive evaluation. Overall, AutoBench-V not only successfully utilizes LVLMs for automated benchmarking but also reveals that LVLMs as judges have significant potential in various domains.
11footnotetext: Equal Contribution. H.B. is the independent researcher.22footnotetext: Corresponding author.33footnotetext: Y.W. and J.Y. are the visiting students at MBZUAI.
@@ 1 Introduction
The flourishing of Large Language Models (LLMs) (Touvron et al., 2023; Achiam et al., 2023; Liu et al., 2024a; Anthropic, 2024) has paved the way for significant advancements in the various downstream applications (Vaswani et al., 2023; Huang et al., 2023; Liu et al., 2023c). As the capabilities of LLMs grew, researchers began to explore the integration of visual information understanding capabilities into LLMs, giving rise to the development of Large Vision-Language models (LVLMs) (Achiam et al., 2023; Liu et al., 2024c). These models are trained on extensive paired image-text datasets, enabling them to perform sophisticated multimodal reasoning by effectively integrating visual and textual information (Zou et al., 2023; Ghandi et al., 2023; Karras et al., 2019; Agrawal et al., 2016; Chen et al., 2024a).
With the widespread adoption of LVLMs, evaluating these models has become increasingly important, for understanding their limitations and reliability better. Recent research (Xu et al., 2023; Liu et al., 2023b; Ying et al., 2024; Li et al., 2023b; a; Yin et al., 2023) emphasize the urgent need for comprehensive and sophisticated evaluation standards that accurately assess LVLMs’ abilities across various modalities. Various benchmarks are aiming to evaluate a range of capabilities of LVLMs including 3D understanding (Yin et al., 2023), perception and cognition capacity (Liu et al., 2023b; Fu et al., 2024), multi-discipline understanding and reasoning (Yue et al., 2024). Even though these works have solidly evaluated certain aspects of LVLMs’ capabilities, they lack the flexibility to support on-demand evaluation across various capability aspects. Recent studies have explored the usage of generative AI in automating evaluation, which offers flexibility in varying evaluation dimensions and reduces the human cost of benchmark dataset construction (Wu et al., 2024; Zhu et al., 2024a; Li et al., 2024). While these studies focus on the automatic evaluation of LLMs, we aim to extend this to visual modality by answering addressing this question: “Can LVLMs serve as a path to automatic benchmarking?”
@@
Automating the evaluation of LVLMs presents several key challenges. First, the targeted capabilities to be evaluated must be clearly identified based on the input demand. This is the foundation that relevant images and appropriate visual question-answering (VQA) tasks can be generated to accurately assess the LVLMs’ performance in those specific aspects. Second, the generated images and VQA tasks should be relevant and accurately reflect the evaluation target. Third, the risk of answer leakage from Examiner LVLM during question generation should be mitigated. This issue arises when the model responsible for generating questions exhibits self-enhancement bias (Ye et al., 2024; Zheng et al., 2023), wherein the model being evaluated is also employed to generate the evaluation cases.
Figure 1: Five key evaluation dimensions supported by AutoBench-V, along with their fine-grained sub-aspects, accompanied by questions and images to assist in understanding.
@@
To address the above challenges, we propose AutoBench-V, which supports automated evaluation of LVLMs based on a user demand regarding specific aspects of model capability (e.g., Spatial Understanding). Initially, the input demand is processed by an examiner LVLM, which categorizes it into several overarching aspects. Each aspect is further divided into several fine-grained components, for which image descriptions of varying difficulty levels are generated. To ensure that the descriptions align with their corresponding images, a self-validation mechanism is applied using VQA (Agrawal et al., 2016). Furthermore, an error control mechanism is implemented to prevent a negative impact on the generation of questions and reference answers. The generated questions and images are then presented to the evaluated LVLM to generate responses, which are assessed against reference answers (Liu et al., 2024b). The pipeline of AutoBench-V is shown in Figure 2.
@@
By leveraging AutoBench-V, we conduct extensive evaluation of seven popular LVLMs across five demanded evaluation capabilities (see Figure 1). The results show that LVLMs exhibit declining performance as task difficulty rises, with varied performances over distinctive LVLMs. While excelling in high-level understanding, they struggle with detailed reasoning, revealing a key area for improvement in future research. We also carried out several human evaluation experiments on the generated cases, which yielded positive results, demonstrating the reliability of our approach. To summarize, our key contributions are three-fold:
@@
$\triangleright$ An automated LVLM evaluation framework. This proposed AutoBench-V is the first automated framework for benchmarking LVLMs’ capability. The framework leverages text-to-image models to generate images for evaluation and employs GPT-4o as an examiner to conduct VQA evaluations. This automation significantly reduces human involvement, enhancing the efficiency and objectivity of the evaluation process.
@@
$\triangleright$ Extensive experiments validating the framework’s effectiveness. We conducted comprehensive experiments, including main evaluations on multiple models, examiner superiority tests, option position bias analysis, and human assessments. The results confirm the framework’s robustness and effectiveness in evaluating LVLMs.
@@
$\triangleright$ In-depth analysis of LVLMs’ performance across diverse visual tasks. Through systematic evaluation with varied user inputs, we find that LVLMs demonstrate strong proficiency in abstract conceptual understanding while exhibiting comparatively lower performance in concrete visual reasoning tasks. These insights offer a perspective on the current state of LVLM technology, highlighting areas with potential for future development and exploration.
Figure 2: An overview of the AutoBench-V pipeline, illustrating the automated evaluation process. It starts from user input intake, then aspect generation, followed by the generation of corresponding images and questions, and finally outputs the evaluation score of LVLMs.
@@ 2 Related Works
Benchmark for LVLMs. The emergence of the LVLMs greatly promoted the development of the multimodal model, demonstrating exceptional progress in their multimodal perception and reasoning capabilities. This makes the past, focused on isolated task performance benchmarks (Karpathy & Fei-Fei, 2015; Agrawal et al., 2016) insufficient to provide a comprehensive evaluation. Subsequent studies have introduced benchmarks for assessing LVLMs across a range of multimodal tasks (Goyal et al., 2017; Lin et al., 2015; Russakovsky et al., 2015). However, these benchmarks often fall short in providing fine-grained assessments of abilities and robust evaluation metrics. Hence, recent works (Liu et al., 2023b; Ying et al., 2024; Fu et al., 2024; Yu et al., 2023; 2024) highlight the critical need for developing advanced, comprehensive benchmarks to more accurately assess LVLMs’ multimodal understanding and reasoning capabilities. However, these benchmarks still have different kinds of limitations. For example, LVLM-eHub (Xu et al., 2023) and LAMM (Yin et al., 2023) have utilized several classical datasets that are widely recognized but not sufficiently novel for current advancements, overlooking the possibility of data leakage during LVLM training. Hence, MMStar (Chen et al., 2024b) aims to solve the unnecessity of visual content and unintentional data leakage that exists in LVLM training via constructing an elite vision-indispensable dataset.
Compared to previous work, AutoBench-V not only automates the entire benchmarking process for LVLMs—significantly reducing human workload and minimizing subjective biases—but also scales up and customizes the evaluation process to address fine-grained user needs.
Automatic benchmarks. The significant early advancements in LLMs have driven the development of various benchmarks designed to automate evaluation processes. For example, LMExamQA (Bai et al., 2023b) employs the concept of a Language-Model-as-an-Examiner to create a comprehensive and scalable evaluation framework. In addition, DyVal (Zhu et al., 2024a) and DyVal2 (Zhu et al., 2024b) both highlight the importance of dynamic assessment, with DyVal focusing on reasoning tasks and DyVal2 adopting a broader psychometric approach. AutoBencher (Li et al., 2024) automates the generation of novel, challenging, and salient datasets for evaluating LLMs, further expanding the scope of automated benchmarking. Other efforts, such as UniGen (Wu et al., 2024) and Task Me Anything (Zhang et al., 2024a), focus on developing more tailored and relevant benchmarks for assessing LLM/LVLMs performance across diverse tasks.
@@
3 AutoBench-V
In this section, we introduce AutoBench-V, a framework designed for automating the process of benchmarking LVLMs, empowered by a LVLM $\mathcal{M}_{v}$ and a text-to-image model $\mathcal{M}_{d}$. As shown in Figure 3, AutoBench-V consists of four modules (Xiao et al., 2024a): user-oriented aspect generation, guided description generation, image generation by self-validation, and test case generation & evaluation.
@@ 3.1 User-Oriented Aspect Generation
User input. The user input can specify an evaluation target focused on certain aspects of LVLMs’ capability. AutoBench-V covers the following key evaluation aspects, which are the most crucial for assessing the capabilities of LVLMs: Basic Understanding, Spatial Understanding (Li et al., 2023b), Semantic Understanding (Meng et al., 2024), Reasoning Capacity (Liu et al., 2023b), and Atmospheric Understanding (Geetha et al., 2024). Notably, the user input is not limited to the above kinds, and can be customized as needed.
@@
3.2 Guided Description Generation
Guidelines formulation. To avoid the generation of irrelevant, abstract, or vague details that could lead to discrepancies in image descriptions, we introduce a guideline generation step (Zhao et al., 2024; Viswanathan et al., 2023). Before generating image descriptions, the LVLM model $\mathcal{M}_{v}$ formulates a guideline $\mathcal{D}_{ij}$ (e.g., in Background vs Foreground aspect, it is essential to distinguish between the elements present in the background and those in the foreground) for each fine-grained aspect $A^{(f)}_{ij}$. This guideline acts as a guideline for $\mathcal{M}_{v}$, ensuring that the generated descriptions are coherent, clear, and specific to the fine-grained aspects under evaluation. The process can be expressed as $\mathcal{D}_{ij}=\mathcal{M}_{v}(A^{(f)}_{ij})$. The generated guideline $\mathcal{D}_{ij}$ is then utilized to guide the subsequent image description $\mathcal{T}_{ij}$.
@@ 3.2 Guided Description Generation
Image description with difficulty grading. To enable a more comprehensive evaluation, we introduce a difficulty-grading mechanism for the image descriptions, which includes the evaluation cases from different difficulties. This is achieved by classifying the generated image descriptions into three difficulty levels: easy, medium, and hard. We show the examples across different difficulties in Figure 4. The difficulty level $d$ is determined by key factors such as background complexity, element relationships, and the intricacy of textures. The generation of $\omega$ image descriptions $\{\mathcal{T}_{ij1}^{d},\mathcal{T}_{ij2}^{d},\dots,\mathcal{T}_{ij\omega}^{d}\}$ for $A^{(f)}_{ij}$ at a specific difficulty level $d$ can be defined as: $\mathcal{\bigcup}_{k=1}^{\omega}\{\mathcal{T}_{ijk}^{d}\}=\mathcal{M}_{v}(q,A^{(f)}_{ij},\mathcal{D}_{ij},d)$, where $d\in\{\text{easy},\text{medium},\text{hard}\}$. This grading strategy allows for a nuanced understanding of the model’s capabilities across a range of challenges with details provided in Appendix C.
@@
Diverse description generation strategy. A key challenge when generating image descriptions at the same difficulty level is minimizing repetitive elements and backgrounds, which can reduce the diversity and generalization of the evaluation. For example, given a user input $q$ related to spatial understanding, the model $\mathcal{M}_{v}$ might tend to produce descriptions centered around urban landscapes, potentially compromising the variety of test cases. To address this, we introduce a description optimization strategy using a semantic graph (Quillian, 1966) to enhance the diversity of image prompts generated by $\mathcal{M}_{v}$, with significant results referred to in Appendix Figure 10. For a visualization of specific words, see Appendix Figure 11 and Figure 12. The process is iterative, and during the $e^{\prime}$th iteration of prompt generation, a topic word $t_{e}$ and a set of $|c|$ related keywords $K_{e}=\{k_{e1},k_{e2},\dots,k_{ec}\}$ are selected. These keywords are added as nodes to the semantic graph $G$, where nodes are connected by edges representing semantic relationships between them.
@@
Formally, let $G_{e}=(V_{e},E_{e})$ be the semantic graph generated at iteration $e$, and let $S_{e}=\left(V_{e-1}\cup\{t_{e}\}\cup K_{e}\right)$. Then $V_{e}=S_{e}\setminus f(S_{e})$ represents the node set of topic words and keywords, and $E_{e}$ is the set of edges capturing the relationships between them. After each round of prompt generation, we apply a degree-based exclusion mechanism, where the number of excluded nodes is determined by a function $f(S_{e})$. This function defines the number of top-degree nodes to be excluded, allowing flexibility in adjusting how many frequently used words are removed as the iterations progress. The function $f(S_{e})$ could be a simple function such as $f(S_{e})=\underset{V^{\prime}\subseteq S_{e},|V^{\prime}|=e}{\text{arg max}}\sum_{v\in V^{\prime}}\deg(v)$, where $\deg(v)$ represents the degree of node $v\in V_{i}$, or it could take a more complex form based on specific conditions. We mitigate redundancy and promote diversity in the generated prompts by excluding these high-degree nodes, which correspond to the most commonly used words. The function $f(S_{e})$ offers the flexibility to control how aggressively the exclusion process operates based on the round number $e$.
@@
$\bigcup_{k=1}^{\omega}\{\mathcal{T}_{ijk}^{de}\}={M}_{v}(q,A^{(f)}_{ij},\mathcal{D}_{ij},V_{e},d),$
where $V_{e}$ represents the refined and diverse set of topic words and keywords after the exclusion mechanism has been applied. We show the detailed procedure in Algorithm Algorithm 1.
@@
0:  User input $q$, model $\mathcal{M}_{v}$, initial set of topic words and keywords $V_{0}$, exclusion function $f(S_{e})$, number of iterations $N$
0:  Set of diverse image descriptions $\{\mathcal{T}(1),\mathcal{T}(2),\dots,\mathcal{T}(N)\}$
1:  Initialize iteration counter $e:=1$
2:  Initialize the set of topic words and keywords $V_{1}:=V_{0}$
3:  while $e\leq N$ do
4:     Select a topic word $t_{e}$ and a set of related keywords $K_{e}=\{k_{e1},k_{e2},\dots,k_{ec}\}$
5:     Form the node set $S_{e}=V_{e-1}\cup\{t_{e}\}\cup K_{e}$
6:     Formulate $E_{e}$, where edge represent semantic relationship.
7:     Identify exclusion set $f(S_{e})=\underset{V^{\prime}\subseteq S_{e},|V^{\prime}|=e}{\text{arg max}}\sum_{v\in V^{\prime}}\deg(v)$
8:     Update the node set as $V_{e}=S_{e}\setminus f(S_{e})$
9:     Set superparameter $d$ and $\omega$
10:     Set $\mathcal{T}(e)=\bigcup_{k=1}^{\omega}\{\mathcal{T}_{ijk}^{de}\}={M}_{v}(q,A^{(f)}_{ij},\mathcal{D}_{ij},V_{e},d)$
11:     Increment the iteration counter: $i:=i+1$
12:  end while
13:  return  Set of diverse image descriptions $\{\mathcal{T}(1),\mathcal{T}(2),\dots,\mathcal{T}(N)\}$

### v3

# AutoBench-V: Can Large Vision-Language Models Benchmark Themselves?

@@ Abstract
Abstract: Large Vision-Language Models (LVLMs) have become essential for advancing the integration of visual and linguistic information. However, the evaluation of LVLMs presents significant challenges as the evaluation benchmark always demands lots of human cost for its construction, and remains static, lacking flexibility once constructed. Even though automatic evaluation has been explored in textual modality, the visual modality remains under-explored. As a result, in this work, we address a question: “Can LVLMs themselves be used to benchmark each other in the visual automatically domain?". We introduce AutoBench-V, an automated framework for serving evaluation on demand, i.e., benchmarking LVLMs based on specific aspects of model capability. AutoBench-V leverages text-to-image models to generate relevant image samples and then utilizes LVLMs to orchestrate visual question-answering (VQA) tasks, completing the evaluation process efficiently and flexibly. Through an extensive evaluation of nine popular LVLMs across five demanded user inputs (i.e., evaluation capabilities), the framework shows effectiveness and reliability.
@@ 1 Introduction
In recent years, the rapid development of natural language processing (NLP) has given rise to Large Language Models (LLMs) (Zhao et al., 2023), which have revolutionized the field and opened new avenues for innovation and practical downstream applications (Liu et al., 2023b, c). As the capabilities of LLMs grow, researchers begin to explore the integration of visual information understanding capabilities into LLMs, leading to the emergence of Large Vision-Language models (LVLMs) (Liu et al., 2024). These models are trained on extensive multi-modal datasets (e.g., image-text data), enabling them to perform sophisticated cross-modal perception, understanding, and reasoning by effectively integrating visual and textual information (Antol et al., 2015; Zou et al., 2023; Ghandi et al., 2023; Chen et al., 2024a).
With LVLMs becoming more widely adopted, their evaluation has become increasingly important for determining their limitations and reliability. Recent research (Xu et al., 2024; Liu et al., 2025; Li et al., 2023b, a; Yin et al., 2024; Ying et al., 2024) highlights the need for evaluation standards that cover a broad spectrum of multimodal capabilities. Although these benchmarks examine selected dimensions of LVLM performance, they lack flexibility for on-demand assessments in different capability areas. Recent studies have explored using generative models in automating evaluation, which offers flexibility in varying evaluation dimensions and reduces the human cost of benchmark dataset construction (Wu et al., 2024; Zhu et al., 2023b; Li et al., 2024). However, most of these efforts concentrate on LLMs operating solely on text, leaving the visual aspect under-explored. Motivated by this gap, we ask the following question: “Can LVLMs themselves be used to automatically benchmark each other in the visual domain?” In other words, is it feasible to leverage LVLMs as both question-generators and evaluators, with minimal human intervention, to create flexible and reliable tests of multimodal understanding?
@@
Automating LVLM evaluation presents several key challenges. First, the targeted capabilities to be evaluated must be identified based on the input demand. This foundation enables the generation of relevant images and appropriate VQA tasks to accurately assess the LVLMs’ performance in those specific aspects [^1] An aspect example would be like “Spatial Understanding” Li et al. (2023b).. Second, the generated content must be relevant and controllable, so that it accurately tests the targeted skills. Third, there is a risk of “answer leakage”: 1) the issue arises when the model responsible for generating questions exhibits self-enhancement bias (Zheng et al., 2023; Ye et al., 2024)—it can achieve higher performance as it embeds too many “hints” in the generated questions, 2) the prior work (Chen et al., 2024b) shows that models can achieve high VQA performance using text alone.
@@
To address the above challenges, we propose AutoBench-V, which supports automated LVLM evaluation based on user demand for specific evaluation aspects of model capability (e.g., spatial understanding). Initially, the input demand (i.e., evaluation aspect) is processed by an examiner LVLM. Each aspect is further divided into several fine-grained components, for which image descriptions of varying difficulty levels are generated. To ensure that descriptions align with their corresponding images, a self-validation mechanism is applied through VQA. Furthermore, an error control mechanism is implemented to prevent a negative impact of misalignment between text and image. To avoid the potential self-enhancement bias and the situation that LVLMs only utilize textual information to guess answers Chen et al. (2024b), we introduce an error-driven option adjustment strategy to reduce such bias and “force” model to utilize visual information for answering. Moreover, we also introduce multi-examiners to avoid being the player and the referee at the same time. Finally, the generated questions and images are then presented to the evaluated LVLM to generate responses, which are assessed by reference answers (Liu et al., 2023a). The high-level pipeline of AutoBench-V is shown in Figure 1.
@@
Using AutoBench-V, we conduct an extensive evaluation of nine popular LVLMs in five evaluation capabilities requested (see Figure 2). The results show that LVLMs exhibit declining performance as task difficulty rises, with varied performances over distinctive LVLMs. While excelling in high-level understanding, they struggle with fine-grained reasoning, revealing a key area for improvement in future research. Moreover, Some models scored below 25% on medium and hard tasks, performing worse than random guessing. To summarize, our key contributions are three-fold:
@@
$\triangleright$ An automated LVLM evaluation framework. The proposed AutoBench-V is the first automated framework for benchmarking LVLMs’ capability. The framework uses text-to-image models to generate images based on textual descriptions, and employs LVLM examiners for VQA evaluations. This automation significantly minimizes human involvement, enhancing efficiency and objectivity in the evaluation process.
@@
$\triangleright$ Extensive experiments to validate the framework’s effectiveness. We conducted comprehensive experiments, including main evaluations on multiple models, ablation studies on each module, option position bias analyses, and human evaluation. These results confirm the framework’s robustness and effectiveness in evaluating LVLMs.
@@
$\triangleright$ In-depth analysis of LVLMs’ performance across diverse visual tasks. Through systematic evaluation with varied user inputs, we obtained many insightful findings. For instance, LVLMs exhibit strong proficiency in abstract conceptual understanding while demonstrating relatively lower performance in concrete visual reasoning tasks. These findings provide insights into the current state of LVLMs, highlighting areas with potential for future development and exploration.
Figure 2: Five key evaluation dimensions in AutoBench-V—Basic Understanding, Reasoning Capacity, Semantic Understanding, Spatial Understanding, and Atmospheric Understanding—and their sub-aspects, illustrated by the representative question–image pairs.
@@ 2 Related Works
Benchmark for LVLMs. The emergence of LVLMs greatly promoted the development of multimodal models, showcasing exceptional progress in their multimodal perception and reasoning capabilities. This makes previous benchmarks, which focused on isolated task performance (Karpathy and Fei-Fei, 2015; Antol et al., 2015) insufficient to provide a comprehensive evaluation. Subsequent studies have introduced various benchmarks to assess LVLMs across a range of multimodal tasks (Goyal et al., 2017; Lin et al., 2014; Russakovsky et al., 2015). However, these benchmarks often lack fine-grained assessments of capabilities and robust evaluation metrics. Hence, recent works (Liu et al., 2025; Ying et al., 2024; Fu et al., 2024; Yu et al., 2023, 2024; Zhou et al., 2024b) underscore the critical need for developing advanced, comprehensive benchmarks to more accurately assess multimodal understanding and reasoning capabilities of LVLMs. However, these benchmarks still have various limitations. For example, LVLM-eHub (Xu et al., 2024) and LAMM (Yin et al., 2024) have utilized several classical datasets that are widely recognized but not sufficiently novel for current advancements, overlooking the possibility of data leakage during LVLM training. Hence, MMStar (Chen et al., 2024b) aims to solve the issure of unnecessary visual content and unintentional data leakage in LVLM training by constructing an elite vision-indispensable dataset.
Automatic benchmarks. The significant early advancements in LLMs have driven the development of various benchmarks designed to automate evaluation processes. For example, LMExamQA (Bai et al., 2024) employs the concept of a Language-Model-as-an-Examiner to create a comprehensive and scalable evaluation framework. In addition, DyVal (Zhu et al., 2023b) and DyVal2 (Zhu et al., 2024) both highlight the importance of dynamic assessment, with DyVal focusing on reasoning tasks and DyVal2 adopting a broader psychometric approach. AutoBencher (Li et al., 2024) automates the generation of novel, challenging, and salient datasets for evaluating LLMs, further expanding the scope of automated benchmarking. Other efforts, such as UniGen (Wu et al., 2024) and Task Me Anything (Zhang et al., 2024a), focus on developing more customized and relevant benchmarks to evaluate LVLM and LLM performance across diverse tasks.
@@
3 The Proposed AutoBench-V Framework
In this section, we introduce AutoBench-V, a framework for automated benchmarking of LVLMs. AutoBench-V makes use of two key components—an LVLM examiner $\mathcal{M}_{v}$ and a text-to-image model $\mathcal{M}_{d}$—to support robust and scalable evaluations. As illustrated in Figure 3, the framework consists of four modules: user-oriented aspect generation, guided description generation, image generation with self-validation, and test-case generation and evaluation. AutoBench-V aims to address three primary challenges:
  • Diverse Evaluation Aspects. To ensure comprehensive benchmarking, AutoBench-V covers various capability dimensions, each graded by difficulty. The breadth is achieved via a hierarchical aspect-guided strategy enriched by semantic graph-based prompt generation.
  • Truthful Evaluation Cases. The framework includes a self-validation step that detects potential errors in generated items. An additional error-control mechanism reinforces factual correctness, improving its reliability.
  • Minimized Evaluation Bias. To reduce bias, AutoBench-V maintains a set of LVLM examiners, limiting self-enhancement bias (Ye et al., 2024) that can occur if a single model handles both question generation and responses. Furthermore, error-driven option adjustment thwarts text-only strategies for guessing answers, thereby increasing the difficulty of evaluation items.
@@ 3.1 User-Oriented Aspect Generation
In this part we detail the implementation of AutoBench-V and for the math notations used below are additionally referred to in the Appendix E.
User input. The user input can specify an evaluation target focused on certain aspects of LVLMs’ capability. AutoBench-V covers the following key evaluation aspects, which are the most crucial for assessing the capabilities of LVLMs: Basic Understanding, Spatial Understanding (Li et al., 2023b), Semantic Understanding (Meng et al., 2024), Reasoning Capacity (Liu et al., 2025), and Atmospheric Understanding (Geetha et al., 2024). Notably, the user input is not limited to the above kinds, and can be customized as needed.
@@
3.2 Diverse Description Generation
@@
Image description with difficulty grading. To enable a more comprehensive evaluation, we introduce a difficulty-grading mechanism for the image descriptions, which includes the evaluation cases from different difficulties. We define three difficulty levels: easy, medium, and hard. We show the examples across different difficulties in Figure 9. The difficulty level $d$ is determined by key factors such as background complexity, element relationships, and the intricacy of textures, which are carefully discussed and designed from the perspective of visual perception. The generation of $\omega$ image descriptions $\{\mathcal{T}_{ij1}^{d},\mathcal{T}_{ij2}^{d},\dots,\mathcal{T}_{ij\omega}^{d}\}$ for $A^{(f)}_{ij}$ at a specific difficulty level $d$ can be defined as: $\mathcal{\bigcup}_{k=1}^{\omega}\{\mathcal{T}_{ijk}^{d}\}=\mathcal{M}_{v}(q,A^{(f)}_{ij},d)$, where $d\in\{\text{easy},\text{medium},\text{hard}\}$. More details about the difficulty grading are provided in Appendix C.
@@
Semantic graph-constraint description generation. A key challenge when generating image descriptions at the same difficulty level is minimizing repetitive elements and backgrounds, which can reduce the diversity and generalization of the evaluation. For example, given a user input $q$ related to spatial understanding, the model $\mathcal{M}_{v}$ might tend to produce descriptions centered around urban landscapes, potentially compromising the variety of test cases. To address this, we introduce a description optimization strategy using a semantic graph (Quillian, 1966) to enhance the diversity of image prompts generated by $\mathcal{M}_{v}$, with significant results referred to Figure 8 in Appendix A. Moreover, we show a visualization of specific words in Figure 10 and Figure 11 in Appendix A. The semantic graph is designed to assist AutoBench-V in avoiding the generation of duplicate descriptions by serving as an iteratively updated structure tailored to prompts within the specified aspect. To illustrate the construction process, without loss of generality. Considering $e^{\prime}$th iteration of prompt generation, a LLM identified topic word $t_{e}$ and a set of $|c|$ related keywords $K_{e}=\{k_{e1},k_{e2},\dots,k_{ec}\}$ are selected. These keywords are added as nodes to the semantic graph $G$, where nodes are connected by edges representing semantic relationships between them.
@@
Formally, let $G_{e}=(V_{e},E_{e})$ be the semantic graph generated at iteration $e$, and let $S_{e}=\left(V_{e-1}\cup\{t_{e}\}\cup K_{e}\right)$. Then $V_{e}=S_{e}\setminus f(S_{e})$ represents the node set of topic words and keywords, and $E_{e}$ is the set of edges capturing the relationships between them. After each round of prompt generation, we apply a degree-based exclusion algorithm, where the number of excluded nodes is determined by a function $f(S_{e})$. This function defines the number of top-degree nodes to be excluded, allowing flexibility in adjusting how many frequently used words are removed as the iterations progress. Here, the function $f(S_{e})$ can be defined by users’ requirements, and in our experiment we define it as $f(S_{e})=\underset{V^{\prime}\subseteq S_{e},|V^{\prime}|=e}{\text{arg max}}\sum_{v\in V^{\prime}}\deg(v)$, where $\deg(v)$ represents the degree of node $v\in V_{i}$, or it could take a more complex form based on specific conditions. We mitigate redundancy and promote diversity in the generated prompts by excluding these high-degree nodes, which correspond to the most commonly used words. The function $f(S_{e})$ offers the flexibility to control how aggressively the exclusion process operates based on the round number $e$.
@@
$\bigcup_{k=1}^{\omega}\{\mathcal{T}_{ijk}^{de}\}={M}_{v}(q,A^{(f)}_{ij},V_{e},d),$
where $V_{e}$ represents the refined and diverse set of topic words and keywords after the exclusion mechanism has been applied. We show the detailed procedure in Algorithm 1.
@@
1:User input $q$, model $\mathcal{M}_{v}$, initial set of topic words and keywords $V_{0}$, exclusion function $f(S_{e})$, number of iterations $N$
2:Set of diverse image descriptions $\{\mathcal{T}(1),\mathcal{T}(2),\dots,\mathcal{T}(N)\}$
3:Initialize iteration counter $e\leftarrow 1$
4:Initialize the set of topic words and keywords $V_{1}\leftarrow V_{0}$
5:while $e\leq N$ do
6:    Select a topic word $t_{e}$ and a set of related keywords $K_{e}=\{k_{e1},k_{e2},\dots,k_{ec}\}$
7:    Form the node set $S_{e}=V_{e-1}\cup\{t_{e}\}\cup K_{e}$
8:    Formulate $E_{e}$, where edges represent semantic relationships.
9:    Identify exclusion set
$f(S_{e})=\underset{V^{\prime}\subseteq S_{e},|V^{\prime}|=e}{\arg\max}\sum_{v\in V^{\prime}}\operatorname{deg}(v)$
10:    Update the node set as $V_{e}\leftarrow S_{e}\setminus f(S_{e})$
11:    Set hyperparameters $d$ and $\omega$
12:    Set
$\mathcal{T}(e)=\bigcup_{k=1}^{\omega}\{\mathcal{T}_{ijk}^{de}\}=\mathcal{M}_{v}(q,A^{(f)}_{ij},\mathcal{D}_{ij},V_{e},d)$
13:    Increment the iteration counter: $e\leftarrow e+1$
14:end whilereturn Set of diverse image descriptions $\{\mathcal{T}(1),\mathcal{T}(2),\dots,\mathcal{T}(N)\}$